Which of the following is not classified as portfolio income for tax purposes? a.Dividend income from stock. b.Interest income on savings accounts. c.Dividends paid from a credit union. d.Net rental income from real estate partnership. e.All of these choices are classified as portfolio income.

John owns a second home in Palm Springs, CA. During the year, he rented the house for $5,000 for 56 days and used the house for 14 days during the summer. The house remained vacant during the remainder of the year. The expenses for the home included $5,000 in mortgage interest, $850 in property taxes, $900 for utilities and maintenance, and $3,500 of depreciation. What is John's deductible rental loss, before considering the passive loss limitations?

a.$875

b.$200

c.$3,200

d.$2,500

e.$0

Susan, a single taxpayer, bought her home 25 years ago for $30,000. She has lived in the home continuously since she purchased it. In 2018, she sells her home for $200,000. What is Susan's taxable gain on the sale?

a.$0

b.$250,000

c.$20,000

d.$170,000

In November 2018, Ben and Betty (married, filing jointly) have a long-term capital gain of $54,000 on the sale of stock. They have no other capital gains and losses for the year. Their ordinary income for the year after the standard deduction is $72,500, making their total taxable income for the year $126,500 ($72,500 + $54,000). What will be their 2018 total tax liability assuming a tax of $8,322 on the $72,500 of ordinary income?

The tax rates on long-term capital gains are as follows:


Income Level
Long-Term
Capital Gains Rates
Married filing jointly
$0–$77,200 0%
$77,201–$479,000 15%
> $479,000 20%


a.$16,422

b.$15,717

c.$8,322

d.$19,712

Vijay sells land and receives $5,000 cash, a motorcycle worth $1,600, and two tickets to the Super Bowl with a total face value (cost) of $800 but worth $1,200. In addition, the buyer assumes the mortgage on the land of $12,000. What is Vijay's amount received in this transaction?

a.$5,000

b.$19,800

c.$7,800

d.$20,200

e.$8,200

Yasmeen purchases stock on January 30, 2017. If she wishes to achieve a long-term holding period, what is the first date that she can sell the stock as a long-term gain?

a.July 31, 2017

b.July 30, 2017

c.February 1, 2018

d.January 20, 2018

e.January 31, 2018
Answer #1

Part 1

Answer is option D

d. Net rental income from real estate partnership

Dividend income and interest income are classified as portfolio income.

Part 2

Answer is option C

c. $3,200

Deduction = 5000-((5000+850+900+3500)*56/70) = $3200

Part 3

Answer is option A

a. $0

Gain = 200000-30000 = $170000

However, there is exclusion up to $250000

Part 4

Answer is option B

b. $15,717

total tax liability = 8322+(49300*15%) = $15717

126500-77200 = 49300

Part 5

Answer is option B

b. $19,800

Amount received = cash + motorcycle + “Worth" of the 2 tickets +land = 5000+1600+1200+12000 = $19800

Part 6

e. January 31, 2018

for long-term gain the asset should be hold for more than one year.